Let A, B, C, D denote, respectively, art, biology, chemistry, and drama courses.


Find the number N of students in a dormitory given the data:


12 take A, 5 takeAand B, 4 takeB and D, 2 take B, C,D,


20 take B, 7 takeAand C, 3 takeC and D, 3 take A, C,D,


20 take C, 4 takeAand D, 3 take A, B,C, 2 take all four,


8 take D, 16 takeB and C, 2 take A, B, D, 71 take none.

1
Expert's answer
2022-01-03T17:01:13-0500

Let T be the number of students who take at least one course. By the Inclusion-Exclusion Principle Theorem, "T=s_{1}-s_{2}+s_{3}-s_{4}" where:

"\begin{array}{ll}

s_{1}=12+20+20+8=60, & s_{2}=5+7+4+16+4+3=39, \\

s_{3}=3+2+2+3=10, & s_{4}=2 .

\end{array}"

Thus T=29, and N=71+T=100.